# JiraiBench: A Bilingual Benchmark for Evaluating Large Language Models’ Detection of Human Self-Destructive Behavior Content in Jirai Community

### v2

@@ 3.2 Data Annotation
Our annotation protocol employs a multidimensional framework to identify self-destructive behaviors in social media texts, establishing three distinct categories: OD, ED, and SH. This breakdown enables finer-grained category-based analysis. For each post, the annotators evaluated the content using a three-point ordinal scale: ”0” indicating absence of targeted behaviors, ”1” signifying first-person expressions, and ”2” denoting third-party descriptions - a distinction that provides critical contextual information for analysis and intervention systems.
The annotation process followed a systematic methodology in which annotators first performed comprehensive readings of each post, including textual content and paralinguistic markers, before identifying potentially harmful content. Annotators were trained to distinguish genuine expressions of harmful tendencies from figurative language, with specific guidelines established for ambiguous cases. We developed precise annotation criteria for each dimension. OD identified substance misuse or dangerous drug consumption; ED captured extreme food restrictions, binge-purge behaviors, or unhealthy weight fixations; and SH encompassed suicidal ideation or self-injury behaviors. Each dimension was evaluated independently, allowing for multi-label classification when appropriate.
For this research, we recruited a team of six annotators, evenly divided between Chinese and Japanese, and spent 100 hours in annotators training and labeling on each side. For each language, the team consisted of two native speakers and one domain expert, resulting in three annotators for Chinese (including two native Chinese speakers and one domain expert) and three annotators for Japanese (including two native Japanese speakers and one domain expert). All annotators received equitable compensation at minimum wage levels applicable in their respective regions (Hong Kong SAR for Chinese and Japan for Japanese). Annotator disagreements were systematically resolved through a combination of majority voting and expert supervision, ensuring annotation consistency. We evaluated inter-annotator reliability using both pairwise (Cohen’s Kappa) and overall (Fleiss’ Kappa) agreement metrics. The results demonstrate substantial agreement across all annotation dimensions, with average Cohen’s Kappa values ranging from 0.69 to 0.78 and Fleiss’ Kappa values consistently above 0.68 for all three categories. These robust agreement scores indicate strong consistency among annotators, confirming the reliability of our annotation framework for the detection of self-destructive behaviors. Specific annotation guidelines can be found in the Appendix B.
@@ 3.3 Data Composition
Our analysis includes two datasets: 10,419 posts from the Chinese social platform Sina Weibo and 5,000 posts from Japanese X (formerly Twitter). The Chinese dataset shows higher content related rates in all categories, with OD-related posts that contain 30. 55% first-person expressions compared to only 3.82% in the Japanese dataset. Similarly, the content of ED in the Chinese data set showed 15.04% first-person expressions versus 3.14% in the Japanese dataset. SH content demonstrated more comparable distributions, with first-person expressions comprising 10.27% of Chinese posts and 8.36% of Japanese posts. Table 1 provides a comprehensive breakdown of both datasets across all behavior categories and label types.
@@
Drug Overdose (OD) Non-concerning (0) 6,268 60.16% 4,706 94.12%
@@
Self-Harming (SH) Non-concerning (0) 8,924 85.65% 4,427 88.54%
@@
4 Experiments
Our experimental framework systematically evaluates the performance of LLMs in detecting self-destructive behaviors in social media content in multiple languages and test conditions. We employ four state-of-the-art language models, Llama-3.1 8B (Meta, 2024), Qwen-2.5 7B (Qwen et al., 2025), DeepSeek-v3 (DeepSeek-AI, 2025), as well as our finetuned JiraiLLM-Qwen. We then structure our investigations around three complementary experimental paradigms.
In our initial experimental design, we attempted to include GPT-4o as a baseline; however, the model consistently refused to perform classification tasks involving self-destructive content despite varied prompting strategies. This limitation highlights significant limitations in the evaluation of closed source models for the detection of sensitive content and underscores the need for specialized research interfaces that balance safety mechanisms with legitimate research on harmful content detection.
4.1 Baseline
The baseline experiments establish fundamental performance benchmarks across both Chinese and Japanese corpora under zero-shot and few-shot learning conditions. For zero-shot evaluation, models receive task instructions without exemplars, requiring them to leverage pre-trained knowledge for classification across our three-dimensional annotation scheme. The few-shot configuration provides two examples representing diverse manifestations of self-destructive behaviors with balanced representation across categories. To avoid selective bias, all examples shown in the data are randomly sampled from our dataset. We independently calculated precision, recall, and F1 scores for each dimension to assess overall classification effectiveness and identify language-specific processing disparities. Prompts are provided in Appendix C.
4.2 Crosslingual Transfer
The examination of the effectiveness of cross-lingual transfer in our study involved fine-tuning Qwen2.5 with 3,000 randomly sampled data points from a Chinese dataset utilizing Chinese prompts, resulting in the development of JiraiLLM-Qwen. Our comprehensive evaluation framework assessed the model’s performance across both the source Chinese dataset and the target Japanese dataset, providing empirical insights into the model’s capacity to transfer linguistic knowledge between related but distinct language systems without explicit training on the target language. This investigation offers valuable insight into the capabilities of cross-linguistic generalization in LLMs, particularly in East Asian language contexts where shared writing systems and linguistic features may facilitate knowledge transfer despite significant structural and lexical differences between languages. Detailed training parameters and hyperparameter configurations are attached in the appendix D.
4.3 Evaluation Metric
To align with the established research norm (Yang et al., 2024), we use the Macro F1 score as the evaluation metrics for our self-destructive behavior detection task. The metric assesses the model’s performance in successfully identifying self-destructive behaviors.
Dataset Prompt Method Task Qwen2.5 7B Llama3.1 8B DeepSeek-v3 Jirai-Qwen Random
Chinese Chinese zero-shot OD 0.5052 0.3598 0.5478 0.6398 0.2504
ED 0.4706 0.3349 0.4163 0.6503 0.3015
SH 0.3927 0.3232 0.4205 0.5533 0.3076
two-shot OD 0.4190 0.4400 - - 0.2504
ED 0.2951 0.3590 - - 0.3015
SH 0.4284 0.3714 - - 0.3076
Japanese zero-shot OD 0.5226 0.3969 0.5229 0.6371 0.2504
ED 0.5431 0.3403 0.3962 0.6069 0.3015
SH 0.4154 0.3240 0.3670 0.5402 0.3076
two-shot OD 0.3078 0.2503 - - 0.2504
ED 0.3470 0.3014 - - 0.3015
SH 0.3118 0.3081 - - 0.3076
English zero-shot OD 0.4796 0.3516 0.5087 0.4769 0.2504
ED 0.4023 0.2771 0.4148 0.6173 0.3015
SH 0.3933 0.3164 0.4286 0.5993 0.3076
two-shot OD 0.3304 0.3352 - - 0.2504
ED 0.2223 0.2173 - - 0.3015
SH 0.4278 0.2773 - - 0.3076
Japanese Chinese zero-shot OD 0.3813 0.2869 0.5521 0.3167 0.3232
ED 0.4139 0.2573 0.4097 0.4163 0.3184
SH 0.4241 0.3280 0.4159 0.3114 0.3131
two-shot OD 0.4087 0.3254 - - 0.3232
ED 0.4172 0.2459 - - 0.3184
SH 0.4319 0.3448 - - 0.3131
Japanese zero-shot OD 0.4059 0.2430 0.5974 0.5117 0.3232
ED 0.4665 0.2000 0.4585 0.5761 0.3184
SH 0.4292 0.2255 0.4782 0.4840 0.3131
two-shot OD 0.3428 0.3232 - - 0.3232
ED 0.3684 0.3184 - - 0.3184
SH 0.3131 0.3130 - - 0.3131
English zero-shot OD 0.3683 0.2385 0.5335 0.4939 0.3232
ED 0.3777 0.2661 0.4282 0.5328 0.3184
SH 0.4302 0.3028 0.4198 0.4738 0.3131
two-shot OD 0.4210 0.2141 - - 0.3232
ED 0.3940 0.2109 - - 0.3184
SH 0.4648 0.2795 - - 0.3131
5 Result and Discussion
Table 2 presents the self-destructive behavior detection outcomes for all models, showing that Jirai-qwen achieves the highest performance on Chinese data across all three detection dimensions. However, both open-source models exhibit a remarkable pattern where Japanese instruction prompts consistently outperform Chinese prompts when processing Chinese content, despite Japanese not being the native language of the content. This indicates a significant cross-cultural transfer effect in self-destructive content detection.
5.1 Effects on Different Instruction Language
Our experimental findings reveal compelling emergent cross-cultural transfer patterns that challenge conventional assumptions about language-model alignment in content moderation tasks. The observed superiority of Japanese instruction prompts when processing Chinese content, especially in a zero-shot scenario, demonstrates a fascinating linguistic-cultural bridge effect that transcends straightforward language matching. This phenomenon manifests consistently across multiple model architectures, suggesting a systematic cross-cultural transfer mechanism rather than model-specific behaviorWendler et al. (2024).
The effectiveness of Japanese instructions for the moderation of Chinese content probably stems from the cultural relationship between these language communities. Given that Jirai cultural frameworks originated in Japan before spreading to Chinese-speaking regions, Japanese prompts appear to activate more nuanced cultural schemata relevant to self-destructive behaviors. These culturally embedded frames potentially guide model attention toward subtle linguistic patterns that Chinese instructions fail to adequately highlight, despite being the native language of the content.
Our findings on cross-cultural transfer patterns in self-destructive content detection align with mechanistic interpretability research by Wendler et al., who identified a three-phase processing pattern in Llama-2 models handling multilingual tasks. Their analysis revealed that latent embeddings in middle processing layers already encode appropriate concepts but consistently assign higher probability to English tokens than to target language equivalents—even when prompted in non-English languages. This computational evidence supports our observation that Japanese instruction prompts outperform Chinese prompts when processing Chinese content. Wendler et al.’s research suggests LLMs process information through an English-biased conceptual layer that enables effective cross-cultural transfer between Japanese and Chinese due to their shared cultural frameworks rather than linguistic similarities. Their dimensional analysis suggests that shared logographic writing systems create additional pathways for cultural concepts to traverse linguistic boundaries—particularly important when detecting nuanced psychological phenomena like self-destructive behaviors where semantic representations may activate more effectively through culturally proximate instruction languages.
This finding has broader implications that could potentially generalize to many similar cross-cultural scenarios where shared cultural history creates linguistic bridges between instruction languages and target behaviors. In domains ranging from mental health screening to content moderation, using instruction languages with historical-cultural connections to target phenomena might enhance detection capabilities beyond what standard native language approaches achieve. This suggests an important methodological consideration for multilingual AI deployment: the optimal instruction language may not be the target dataset language, but rather one that activates relevant cultural schema for the specific detection task in question.
5.2 Effects on Zero-shot VS Few-shot prompting
The contrasting performance patterns between zero-shot and few-shot methodologies in our experimental framework reveal nuanced dynamics in the detection of cross-cultural self-destructive behavior. Our analysis demonstrates that zero-shot approaches consistently outperform few-shot methods across most model language configurations, contradicting conventional wisdom regarding few-shot learning benefits. This phenomenon likely stems from the culturally embedded nature of self-destructive behaviors, which manifest through subtle linguistic and contextual patterns that resist straightforward exemplification. The cultural scheme underlying self-destructive behaviors in East Asian contexts, particularly within Jirai cultural frameworks originating from Japan, involve deeply situated knowledge that few-shot exemplars struggle to adequately represent. These cultural understandings encompass complex sociolinguistic signals, indirect expressions, and culturally specific metaphors that may be entirely absent from model training distributions.
The performance degradation observed in few-shot settings potentially indicates interference effects, whereby provided examples inadvertently constrain model attention to narrow manifestations of self-destructive ideation that fail to generalize across the diverse expression patterns present in authentic data (Yang et al., 2022). This limitation appears particularly pronounced when the few-shot examples lack the cultural depth necessary to activate appropriate interpretive frameworks. The complex interrelationship between instruction language and exemplar effectiveness further compounds these challenges, suggesting that the cultural alignment between prompt language and dataset requires more sophisticated calibration than few-shot learning readily accommodates.
6 Conclusion
JiraiBench represents the first bilingual benchmark for evaluating LLMs’ capability to detect self-destructive content across Chinese and Japanese online communities. Our comprehensive experiments across four state-of-the-art models reveal significant limitations in current systems’ effectiveness in identifying harmful content within these linguistically and culturally complex environments. The observed emergent pattern—where Japanese instruction prompts consistently outperform Chinese prompts for Chinese content detection—underscores critical cultural-linguistic alignment effects stemming from the Jirai phenomenon’s Japanese origins. This counterintuitive discovery demonstrates that cultural proximity can sometimes outweigh linguistic similarity in cross-lingual detection tasks.
Future work should focus on developing more robust cross-cultural transfer learning methodologies, expanding benchmark datasets to include additional languages and cultural contexts, and incorporating more nuanced annotation schemes that capture the complex manifestations of self-destructive behaviors across diverse communities.
Ethical Statement
This research addresses the complex ethical terrain surrounding self-destructive behavior detection in multilingual online communities. Our work leverages publicly accessible social media data containing sensitive mental health content, including drug overdose , eating disorders , and self-harm, necessitating rigorous ethical considerations throughout our methodological framework.
We acknowledge inherent sampling limitations in our approach, as user-generated content on social media platforms reflects demographic, linguistic, and cultural disparities that may affect the generalizability of our findings. The Jirai community itself represents a specific transnational online subculture whose linguistic patterns and behavioral expressions may differ from those in other vulnerable populations or platforms. These constraints are transparently documented to contextualize our results appropriately.
Our data collection and annotation protocols prioritized ethical treatment of both content and annotators. All six annotators received comprehensive training in recognizing harmful content patterns while maintaining appropriate emotional distance. To reflect the emotionally intensive nature of content evaluation and the specialized expertise required, all annotators received compensation substantially exceeding minimum wage standards in their respective regions (Hong Kong SAR for Chinese and Japan for Japanese).
We implemented stringent privacy preservation measures in strict adherence to platform policies. All datasets underwent comprehensive desensitization procedures to eliminate personally identifiable information, including usernames, timestamps, location data, and other potential identifiers. Given the ethical considerations surrounding informed consent, protection of user privacy, and potential stigmatization of the studied community, we have implemented a controlled data release policy wherein both language datasets will be available exclusively upon request to verified academic and research institutions.
Our research contributes primarily to enhancing content moderation systems that increasingly incorporate AI technologies. By examining cross-lingual and cross-cultural dimensions of self-destructive behavior detection, we aim to develop more culturally informed approaches that can better support vulnerable individuals across linguistic boundaries. This work aligns with broader ethical imperatives in the NLP community to develop responsible AI systems that can identify concerning content patterns while respecting principles of privacy, cultural sensitivity, and appropriate intervention methodologies.
Appendix A Data Example

### v3

@@ 3.2 Data Annotation
Our annotation protocol employs a multidimensional framework to identify RHBs in social media texts, establishing three distinct categories: Drug Poisoning (DM), Eating Disorders (ED), and Non-suicidal Self-Injury (NSSI). This breakdown enables fine-grained, category-based analysis. For each post, annotators evaluated the content using a three-point ordinal scale: "0" for the absence of targeted behaviors, "1" for first-person expressions (e.g. "I relapsed again"), and "2" for third-party descriptions or general discussions (e.g. "It is dangerous to mix these pills"), a distinction that provides critical context for analysis and intervention.
The annotation process followed a structured, multi-step methodology. Instead of relying on general impressions, each annotator was required to perform a complete reading, which meant analyzing the full textual content of the post, including paralinguistic markers such as emojis and relevant hashtags that could indicate tone or intent. Following this initial reading, they applied a set of precise annotation criteria, which were developed prior to the main annotation phase to ensure consistency. These criteria, detailed in Appendix C, provided explicit definitions for each category:
  • DM was defined as any clear mention of substance or drug misuse or dangerous levels of drug consumption.
  • ED was defined as content capturing extreme food restrictions, binge-purge behaviors, or an unhealthy fixation on weight.
  • NSSI was defined as discussions of suicidal ideation or direct mentions of self-injury behaviors.
Annotators were specifically trained with these guidelines to distinguish genuine expressions of harmful tendencies from figurative or metaphorical language. Each of the three dimensions (DM, ED, NSSI) was evaluated independently for every post, allowing for multi-label classification where a single post might contain evidence of multiple behaviors.
For this research, we recruited a team of six annotators, evenly divided between Chinese and Japanese, and spent over 100 hours training and labeling the annotators on each side. For each language, the annotation team comprised three members: two native speakers and one community expert, applied to both Chinese and Japanese. All annotators received equitable compensation at minimum wage levels applicable in their respective regions (Hong Kong SAR for Chinese and Japan for Japanese). Annotator disagreements were systematically resolved through a combination of majority voting and expert supervision, ensuring annotation consistency. We evaluated inter-annotator reliability using both pairwise and overall agreement metrics shown in Appendix D. The results demonstrate substantial agreement across all annotation dimensions, confirming the reliability of our annotation framework for the detection of RHBs. Specific annotation guidelines can be found in the Appendix C.
@@ 3.3 Data Composition
Our analysis includes two datasets: 10,419 posts from the Chinese social platform Sina Weibo and 5,000 posts from Japanese X (formerly Twitter). The Chinese dataset shows higher content related rates in all categories, with DM-related posts that contain 30.55% first-person expressions compared to only 3.82% in the Japanese dataset. Similarly, the content of ED in the Chinese data set showed 15.04% first-person expressions versus 3.14% in the Japanese dataset. NSSI content demonstrated more comparable distributions, with first-person expressions comprising 10.27% of Chinese posts and 8.36% of Japanese posts. Table A provides a comprehensive breakdown of both datasets across all behavior categories and label types.
4 Experiments
Our experimental framework systematically evaluates the performance of LLMs in detecting risky health behaviors in social media content in multiple languages and test conditions. We employ four state-of-the-art language models, Llama-3.1 8B (Meta, 2024), Qwen-2.5 7B (Qwen et al., 2025), DeepSeek-v3 (DeepSeek-AI, 2025), as well as our finetuned JiraiLLM-Qwen. We then structure our investigations around three complementary experimental paradigms.
In our initial experimental design, we attempted to include GPT-4o as a baseline; however, the model consistently refused to perform classification tasks that involve the content of the RHB despite various prompting strategies. This limitation highlights significant limitations in the evaluation of closed source models for the detection of sensitive content and underscores the need for specialized research interfaces that balance safety mechanisms with legitimate research on harmful content detection.
4.1 Baseline
The baseline experiments establish fundamental performance benchmarks across both Chinese and Japanese corpora under zero-shot and few-shot learning conditions. For zero-shot evaluation, models receive task instructions without exemplars, requiring them to leverage pre-trained knowledge for classification across our three-dimensional annotation scheme. The few-shot configuration provides two examples that represent diverse manifestations of risky health behaviors with balanced representation across categories. To avoid selective bias, all examples shown in the data are randomly sampled from our dataset. We independently calculated precision, recall, and F1 scores for each dimension to assess overall classification effectiveness and identify language-specific processing disparities. The prompts are provided in the appendix E.
4.2 Crosslingual Transfer
The evaluation of the effectiveness of cross-lingual transfer in our study involved fine-tuning Qwen2.5 with 3,000 randomly sampled data points from a Chinese data set utilizing Chinese prompts, resulting in the development of JiraiLLM-Qwen. Our comprehensive evaluation framework assessed the model’s performance across both the source Chinese dataset and the target Japanese dataset, providing empirical insights into the model’s capacity to transfer linguistic knowledge between related but distinct language systems without explicit training on the target language.
This investigation offers valuable insight into the capabilities of cross-linguistic generalization in LLMs, particularly in East Asian language contexts where shared writing systems and linguistic features may facilitate knowledge transfer despite significant structural and lexical differences between languages. Detailed training parameters and hyperparameter configurations are attached in the appendix F.
4.3 Evaluation Metric
To align with the established research norm (Yang et al., 2024), we use the Macro F1 score as the evaluation metrics for our risky health behavior detection task. The metric assesses the model’s performance in successfully identifying risky health behaviors.
Dataset Prompt Method Task Qwen2.5 7B Llama3.1 8B DeepSeek-v3 Jirai-Qwen (CN) Jirai-Qwen (JP) Random
Chinese Chinese zero-shot DM 0.5052 0.3598 0.5478 0.692 0.6914 0.2504
ED 0.4706 0.3349 0.4163 0.6503 0.6997 0.3015
NSSI 0.3927 0.3232 0.4205 0.573 0.5714 0.3076
two-shot DM 0.4190 0.4400 - - - 0.2504
ED 0.2951 0.3590 - - - 0.3015
NSSI 0.4284 0.3714 - - - 0.3076
Japanese zero-shot DM 0.5226 0.3969 0.5229 0.6371 0.6557 0.2504
ED 0.5431 0.3403 0.3962 0.6069 0.6685 0.3015
NSSI 0.4154 0.3240 0.3670 0.5402 0.5761 0.3076
two-shot DM 0.3078 0.2503 - - - 0.2504
ED 0.3470 0.3014 - - - 0.3015
NSSI 0.3118 0.3081 - - - 0.3076
English zero-shot DM 0.4796 0.3516 0.5087 0.4769 0.5600 0.2504
ED 0.4023 0.2771 0.4148 0.6173 0.5912 0.3015
NSSI 0.3933 0.3164 0.4286 0.5993 0.5704 0.3076
two-shot DM 0.3304 0.3352 - - - 0.2504
ED 0.2223 0.2173 - - - 0.3015
NSSI 0.4278 0.2773 - - - 0.3076
Japanese Chinese zero-shot DM 0.3813 0.2869 0.5521 0.3167 0.5007 0.3232
ED 0.4139 0.2573 0.4097 0.4163 0.4514 0.3184
NSSI 0.4241 0.3280 0.4159 0.3114 0.3600 0.3131
two-shot DM 0.4087 0.3254 - - - 0.3232
ED 0.4172 0.2459 - - - 0.3184
NSSI 0.4319 0.3448 - - - 0.3131
Japanese zero-shot DM 0.4059 0.2430 0.5974 0.5117 0.4366 0.3232
ED 0.4665 0.2000 0.4585 0.5761 0.4491 0.3184
NSSI 0.4292 0.2255 0.4782 0.4840 0.3800 0.3131
two-shot DM 0.3428 0.3232 - - - 0.3232
ED 0.3684 0.3184 - - - 0.3184
NSSI 0.3131 0.3130 - - - 0.3131
English zero-shot DM 0.3683 0.2385 0.5335 0.4939 0.4106 0.3232
ED 0.3777 0.2661 0.4282 0.5328 0.4151 0.3184
NSSI 0.4302 0.3028 0.4198 0.4738 0.3844 0.3131
two-shot DM 0.4210 0.2141 - - - 0.3232
ED 0.3940 0.2109 - - - 0.3184
NSSI 0.4648 0.2795 - - - 0.3131
5 Result and Discussion
Table 1 presents the RHB detection outcomes for all models, showing that Jirai-qwen achieves the highest performance on Chinese data across all three detection dimensions. However, both open-source models exhibit a remarkable pattern in which Japanese instruction prompts consistently outperform Chinese prompts when processing Chinese content, despite Japanese not being the native language of the content. This indicates a significant cross-cultural transfer effect in RHB content detection.
5.1 Effects on Different Instruction Language
Our experimental findings reveal compelling emergent cross-cultural transfer patterns that challenge conventional assumptions about language-model alignment in content moderation tasks. In particular, we observe that Japanese instruction prompts often outperform Chinese prompts when moderating Chinese content, particularly in zero-shot scenarios. This surprising pattern suggests a linguistic-cultural bridge effect that transcends simple language matching. The phenomenon manifests itself across multiple model architectures, pointing toward a systematic cross-cultural transfer mechanism rather than an artifact of any single model Wendler et al. (2024).
We hypothesize that the effectiveness of Japanese instructions stems from the deep cultural relationship between Japanese and Chinese-speaking communities, particularly in the domain of RHBs, where the "Jirai" cultural framework originated in Japan before being distributed in Chinese-speaking contexts. Japanese prompts may more effectively activate nuanced cultural schemata associated with RHB discourse, directing model attention toward subtle linguistic markers that Chinese instructions, despite their native language advantage, fail to adequately emphasize.
However, this cross-cultural prompt effect shows significant architecture sensitivity, with Japanese prompts producing inconsistent gains across different LLMs. Models like Qwen2.5 7B can incorporate region-specific pretraining data or tokenization schemes that favor prompt-content language alignment. Safety tuning and censorship mechanisms emerge as critical confounding factors, suppressing culturally sensitive tokens differentially across models and impacting generalization. This was particularly evident in the GPT-4o trials, which frequently refused classification tasks involving sensitive content, demonstrating how alignment and safety tuning can inadvertently impede legitimate content moderation research.
Our findings resonate with Wendler et al.’s mechanistic interpretability work, which identified a three-phase processing pattern in Llama-2 models handling multilingual input. Their analysis shows that while middle-layer embeddings successfully encode relevant concepts, they systematically favor English tokens even under non-English prompting, revealing that LLMs navigate through an inherently English-biased conceptual space. This bias paradoxically facilitates effective transfer between culturally proximate languages like Japanese and Chinese, where shared cultural substrates (logographic writing systems, historical conceptual exchange) establish pathways for detecting semantically complex phenomena like RHBs more robustly when instruction languages resonate culturally rather than linguistically with target content. These insights have profound implications for multilingual AI in mental health screening and content moderation, suggesting that optimal instruction language selection depends less on linguistic correspondence and more on cultural schema activation. Strategic instruction selection based on shared cultural histories and conceptual frameworks can unlock superior performance, particularly for culturally embedded and psychologically nuanced content.
Model (train) Eval Prompt OD ED SH Avg
Jirai-Qwen-cn-LoRA CN Chinese 0.6082 0.6000 0.5678 0.5920
Japanese 0.6622 0.6532 0.5743 0.6299
English 0.5476 0.5624 0.5382 0.5494
JP Chinese 0.5071 0.4836 0.3845 0.4584
Japanese 0.4647 0.5120 0.3980 0.4582
English 0.4220 0.4109 0.4062 0.4131
Jirai-Qwen-jp-LoRA CN Chinese 0.6089 0.6001 0.5671 0.5920
Japanese 0.6598 0.6535 0.5735 0.6289
English 0.5475 0.5639 0.5392 0.5502
JP Chinese 0.5068 0.4849 0.3846 0.4587
Japanese 0.4721 0.5100 0.3962 0.4595
English 0.4218 0.4119 0.4064 0.4134
Table 2: F1 scores by evaluation language, prompt language, and task for two Qwen2.5-7B LoRA variants. Values are shown to 4 decimals.
5.2 Effects on Zero-shot VS Few-shot prompting
The contrasting performance patterns between zero-shot and few-shot methodologies in our experimental framework reveal nuanced dynamics in the detection of cross-cultural risky health behaviors. Our analysis demonstrates that zero-shot approaches consistently outperform few-shot methods across most model language configurations, contradicting conventional wisdom regarding few-shot learning benefits. This phenomenon likely stems from the culturally embedded nature of risky health behaviors, which manifest through subtle linguistic and contextual patterns that resist straightforward exemplification. The cultural scheme underlying RHB in East Asian contexts, particularly within Jirai cultural frameworks originating from Japan, involve deeply situated knowledge that few-shot exemplars struggle to adequately represent. These cultural understandings encompass complex sociolinguistic signals, indirect expressions, and culturally specific metaphors that may be entirely absent from model training distributions.
Performance degradation observed in few-shot settings potentially indicates interference effects, whereby provided examples inadvertently constrain model attention to narrow manifestations of concerning behavioral patterns that fail to generalize across the diverse expression patterns present in authentic data (Yang et al., 2022). This limitation appears particularly pronounced when the few-shot examples lack the cultural depth necessary to activate appropriate interpretive frameworks. The complex interrelationship between instruction language and exemplar effectiveness further compounds these challenges, suggesting that the cultural alignment between prompt language and dataset requires more sophisticated calibration than few-shot learning easily accommodates.
5.3 Cultural vs. Orthographic Transfer
To address the question of whether the advantages of Japanese prompts come from cultural framing or mere script overlap, we conducted an ablation experiment where all Kanji characters were stripped from Japanese prompts and replaced with phonetic kana. This manipulation allows us to isolate the contribution of shared logographic elements while preserving cultural semantic framing. The model’s performance remained relatively strong despite the increased difficulty due to the loss of semantic precision in Kanji-to-Hiragana mapping.
Table 3 presents the Macro-F1 scores for both the base model (zero-shot) and fine-tuned model across the three detection dimensions. All results are statistically significant using paired t-test ($p<0.05$).
[…]
DM 0.4003 0.4720 +0.0717
ED 0.4220 0.4930 +0.0710
NSSI 0.3587 0.3732 +0.0145
Table 3: Ablation results showing performance with Kanji-stripped Japanese prompts
This result supports our hypothesis that cross-lingual transfer is driven by cultural framing rather than mere script overlap, as performance remains stable even when shared characters are removed. The continued advantage of Japanese prompts,even in a reduced-script setting,demonstrates that semantic framing rooted in the Jirai cultural context plays a larger role than orthographic similarity.
5.4 Fine-tuning for Cross-lingual Transfer
We study whether parameter-efficient fine-tuning improves risky health behavior detection and whether gains transfer across languages, and we compare these findings to our prior fine-tuning experiment (Table 1). Starting from Qwen2.5-7B-Instruct, we train two LoRA variants using the same recipe, JiraiQwen-CN-LoRA and JiraiQwen-JP-LoRA. In our main run, we fine-tune on 3,000 randomly sampled Chinese posts with supervised labels under the Chinese prompt template for 3 epochs on 4 NVIDIA A6000 GPUs (per-device batch size 5, gradient accumulation 10, effective batch size 50), keeping other hyperparameters at default values. We then evaluate on both Chinese and Japanese test sets while varying instruction language (Chinese, Japanese, English) to quantify prompt effects.
Compared to the results of the base model in Table 1, supervised adaptation again produces large improvements in the domain in Chinese. Under Chinese prompts on the Chinese dataset, Qwen2.5 increases from an average Macro-F1 of 0.456 (DM, ED, NSSI) to 0.638 with JiraiQwen (CN), confirming that fine-tuning substantially strengthens detection in the source domain. The JiraiQwen-{CN,JP}-LoRA experiment isolates two factors not previously disentangled.
First, the LoRA variant label has negligible impact: across all evaluation settings, JiraiQwen-CN-LoRA and JiraiQwen-JP-LoRA differ by at most 0.0012 Macro-F1, suggesting that the observed behaviors are not driven by which LoRA variant is used under our recipe.
Second, the effects of the instruction language are strong and asymmetric. In the Chinese test set, the Japanese prompts remain best after fine-tuning (0.629 Macro-F1) compared to Chinese (0.592) and English (0.550), reproducing the advantage of the cross-cultural prompt under controlled adaptation. On the Japanese test set, overall performance remains substantially lower (0.444 averaged across prompts and LoRA variants), Chinese and Japanese prompts are effectively tied (both around 0.459), and English prompts are consistently worse (0.413).
Together, these results show that fine-tuning yields strong in-domain performance, and that instruction language can materially shape what the model attends to at test time. In particular, for Chinese evaluation, Japanese prompts consistently achieve the highest F1 across both CN-LoRA and JP-LoRA, even outperforming Chinese prompts, which is consistent with our hypothesis that Japanese instructions more effectively activate the Jirai-origin cultural schema that structures RHB discourse and later diffused into Chinese-speaking contexts. s.
6 Conclusion
JiraiBench represents the first cross-lingual benchmark for evaluating LLMs’ capability to detect risky health behaviors across Chinese and Japanese online communities. Our comprehensive experiments across four state-of-the-art models reveal significant limitations in current systems’ effectiveness in identifying RHB content within these linguistically and culturally complex environments. The observed emergent pattern, where Japanese instruction prompts consistently outperform Chinese counterparts, underscores critical cultural-linguistic alignment effects stemming from the Jirai phenomenon’s Japanese origins. This counterintuitive discovery demonstrates that cultural proximity can sometimes outweigh linguistic similarity in cross-lingual detection tasks.
Future work should focus on developing more robust cross-cultural transfer learning methodologies, expanding benchmark datasets to include additional languages and cultural contexts, and incorporating more nuanced annotation schemes that capture the complex manifestations of RHBs across diverse communities.
Limitations and Ethical Statement
This research engages with the complex ethical considerations involved in detecting risky health behaviors within multilingual online communities. Given the sensitive nature of the content, such as discussions of drug misuse, eating disorders, and nonsuicidal self-injury, we recognized the importance of ethical oversight. The IRB office at the main institution have confirmed in an official communication that social media research of this nature does not qualify as human subjects research and therefore does not require IRB review or exemption. Accordingly, IRB review was not sought. Our study follows established ethical standards for working with publicly available online data.
We recognize valid concerns about the sensitive nature of these data and the vulnerability of the individuals whose content comprises our data set. The decision to conduct this research was not taken lightly. We believe that the technical contributions, specifically the development of culturally informed detection systems that can identify at-risk individuals across linguistic boundaries, provide significant potential benefits that justify the careful use of these data. Enhanced detection capabilities can enable earlier intervention and support for vulnerable individuals in online communities where traditional mental health resources may be inaccessible.
To address privacy concerns, we implemented comprehensive data anonymization protocols that exceed standard practices.
  • Complete removal of all personally identifiable information (PII), including usernames, user IDs, timestamps, location data, and platform-specific identifiers
  • Systematic review of content to redact any self-disclosed personal information embedded within posts
  • Restriction of raw data access to verified academic institutions with demonstrated ethical review processes
To ensure responsible data handling across our multiinstitutional collaboration, we implemented strict protocols governing author access to sensitive data. All data annotation activities were conducted exclusively on datasets from which personally identifiable information (PII) had been thoroughly removed prior to any collaborative work. Authors affiliated with institutions other than [main institution] worked solely with these anonymized datasets, with several coauthors serving as community experts who contributed specialized knowledge to the annotation process after PII removal. These collaborating authors, including community experts, provided valuable contributions through annotation of deidentified content, model design, methodological discussions, and manuscript preparation without ever accessing raw data containing personal information.
Throughout the research process, all coauthors from external institutions had no direct contact with private or identifiable information, instead working with sanitized datasets while contributing feedback on analytical approaches, sharing code for computational analysis, and providing domain expertise in annotation tasks. This collaborative structure allowed us to benefit from diverse expertise, including specialized community knowledge essential for accurate annotation, while maintaining complete separation from original sensitive datasets and upholding the highest standards of data protection and ethical responsibility in handling sensitive health-related online content.
We acknowledge that despite these measures, the absence of explicit consent from content creators remains an ethical limitation. However, we believe that the potential harm from not developing effective detection systems, which could result in missed opportunities for intervention, outweighs the carefully mitigated risks of our approach. Our controlled data release policy ensures that access is limited to researchers who demonstrate both technical competence and ethical commitment to protecting vulnerable populations.
Our annotation protocols prioritized the well-being of our research team. The six annotators received comprehensive training on the recognition of harmful content patterns while maintaining an appropriate emotional distance, with access to mental health resources throughout the annotation process. Compensation substantially exceeded minimum wage standards in their respective regions, reflecting both the specialized expertise required and the emotionally intensive nature of the work.
We remain committed to the responsible development of AI systems that can identify content patterns concerning while respecting the principles of privacy, cultural sensitivity, and human dignity. Future iterations of this work will explore alternative methodologies that can further minimize privacy risks while maintaining detection effectiveness, including synthetic data generation and federated learning approaches that eliminate the need for centralized data collection.
Acknowledgment
This work was supported by JST CREST Grant Number JPMJCR21M2. This work was also supported by JST ACT-X (Grant JPMJAX24CU) and JSPS KAKENHI (Grant 24K20832).
Appendix A Data Distribution
We present the JiraiBench data distribution in Table 4.
@@ Appendix A Data Distribution
Drug Misuse (DM) Non-concerning (0) 6,268 60.16% 4,706 94.12%
@@
Non-Suicidal Self-Injury (NSSI) Non-concerning (0) 8,924 85.65% 4,427 88.54%
@@
Appendix B Data Example